# BOTH2Hands: Inferring 3D Hands from Both Text Prompts and Body Dynamics

### 20 Dec 2023 v3

@@ Abstract
The recently emerging text-to-motion advances have spired numerous attempts for convenient and interactive human motion generation. Yet, existing methods are largely limited to generating body motions only without considering the rich two-hand motions, let alone handling various conditions like body dynamics or texts. To break the data bottleneck, we propose BOTH57M, a novel multi-modal dataset for two-hand motion generation. Our dataset includes accurate motion tracking for the human body and hands and provides pair-wised finger-level hand annotations and body descriptions. We further provide a strong baseline method, BOTH2Hands, for the novel task: generating vivid two-hand motions from both implicit body dynamics and explicit text prompts. We first warm up two parallel body-to-hand and text-to-hand diffusion models and then utilize the cross-attention transformer for motion blending. Extensive experiments and cross-validations demonstrate the effectiveness of our approach and dataset for generating convincing two-hand motions from the hybrid body-and-textual conditions. Our dataset and code will be disseminated to the community for future research.
* Corresponding author.{strip}
@@ 1 Introduction
The recent text-2-motion advances [56, 25, 8, 67] mostly focus on generating body motions only. In contrast, the convenient generation of two-hand motions from text prompts has significantly fallen behind, mainly due to severe data scarcity. The wider adopted text-motion datasets [32, 13] embrace limited two-hand motions and corresponding textual annotations. Only recently, the concurrent work Motion-X [26] provides a large-scale dataset with expressive human motions and paired text prompts. Yet, it still lacks detailed annotations for the hand motions, making the fine-grained generation challenging, let alone enabling explicit finger-level controls. On the other hand, various methods [37, 45, 44] synthesize two-hand motions with body motion as extra conditions. Such a body-to-hand setting implicitly reasons the inherent correlations of human motions between the body and hands, and hence effectively handles specific scenarios like speeches or daily conversations [73, 64]. However, only body-level reasoning falls short of providing explicit and direct controls of hand motions, especially in a human-interpretable manner like text prompts.
@@
Table 1: Dataset comparisons. We conduct a comparison of datasets that encompass body and hand motions. Vocab. denotes the distinct vocabulary numbers used for annotation. Annotation. refers to text annotations. Hand Joint Standard Deviation. reflects the standard deviation of hand joint positions, indicating the diversity of hand motions. Detailed annotation. refers to the number of text annotations for specific skeleton parts in each motion clip.
@@ 2 Related Works
Motion Generation. Currently, numerous works are focusing on motion generation under various conditions such as text [56, 67, 68, 41, 40, 13, 9, 1], speech [63, 29, 73, 62, 14], music [72, 61], objects [53, 4, 7, 15] and action label [70, 38, 3]. Among these, text-to-motion generation is a popular but challenging task due to the ambiguity of natural language and the difficulty of aligning time and space domain [39, 21]. MotionClip [55] presents a method that aligns text with other modalities, enhancing model mapping text descriptions to motion sequences. As diffusion model [50, 18, 51, 10] was introduced and used in various generative tasks [46, 49, 8, 46, 48], it also performs well in motion generation. Human Motion Diffusion Model [56] combined the diffusion model with text condition and give out a good result. Other motion synthesis works like T2M-GPT [67] applied the transformer architectures in motion generation and proved their effectiveness. MLD (motion-latent-diffusion) [6] made the motion synthesis in latent space. There also exists some works focusing on human interaction scenes like InterGen [25] achieving relatively good results. Full-body motion generation holds considerable significance in some specific domains like human object interaction (HOI) [53, 12, 58, 23, 54] and speech [63, 2, 30, 60], due to its extensive applicability.
@@
Hands are equally important as bodies in motion generation [36], presenting unique challenges due to their high density in smaller spatial occupancy. Previous hand generation works concentrate on physics-based issues [42, 27, 71] such as contact surface deformation when interacting with objects [65], or building realistic parametric models [47, 24]. Other existing data-driven [19, 33, 35, 52] works frequently focus on specific scenarios [66], modeling the individual as a whole unit rather than parts during interaction [11, 59], or are like Body to hands [37], introducing body condition as input and achieving impressive results on hand motion synthesis [45, 44, 64, 73]. Previous studies often focus only on full-body synthesis or body motions only, paying little attention to generating hands aligned with both body motions and text controls.
@@
Motion Dataset. These days various motion datasets have been presented. Some action-labeled datasets like BABEL [43] offer basic control signal alignment with motions, but differ from natural language due to their structure of verb-object phrases, lacking the components found in complete sentences. Datasets such as KIT [40] or HumanML3D [13] provide detailed natural annotations on human motions, are useful for motion generation [56, 67, 6], but do not contain hands, which also plays an important role on human motions. As parametric models of hands [47, 24] are proposed, it is easier to obtain hand data and construct datasets. There exist some datasets [34] with abundant data related to hands and their interactions. Due to technical limitations, they hardly contain both hand and body data. Datasets like GRAB [53] contain rich hand gestures but are mainly narrowed down to HOI scenes. BEAT [28] differs by using speech text aligned with gestures during a speech, deviating from standard motion description. Currently largest full-motion dataset Motion-X [26] contains annotations aligned with motion but lacks finger-level annotations focusing on hands, and the diversity of hand movement is not as rich as their body motions. Since the current 3D full-motion dataset focuses few on daily hands [53, 69, 28, 26], more data with rich daily hand gestures with detailed annotation is needed for body hand motion synthesis.
@@ 3 BOTH57M Dataset
Overview. We introduce the BOTH57M, a unique body-hand motion dataset comprising 1,384 motion clips and 57.4M frames, with 23,477 manually annotated motions and a rich vocabulary of 4,140 words, focusing on hands and body motion in daily various activities, referencing the book “Dictionary of Gestures ” [5] and supplementing with custom-designed movements. To the best of our knowledge, this is the only dataset that provides hybrid and detailed annotations of both body and hands at present, providing infinite possibilities for future tasks. Tab. 1 shows a detailed comparison of various body hand datasets with ours. The rich vocabulary and hand diversity underscores our advantage in tackling the text-body to hand task. For a detailed explanation of our motions collected and body annotations, please refer to supplementary materials.
Data Collecting. We utilize 32 RGB cameras for dense-view at the resolution of 3840$\times$2160 and a frame rate of 59.97 fps, offering sufficient appearance supervision for body-hand motion capturing. All cameras are well calibrated for dealing with occlusion. During data collection, participants are instructed to perform movements listed in the “Dictionary of Gestures ”excluding unfriendly gestures. Volunteers are required to expand the pose in the book into a body-hand motion sequence and perform it. We motivate performers to diversify hand gestures and body motions while maintaining the original meaning of the book. Subsequently, manual annotation is implemented. Each motion is segmented into 5-second clips and annotated independently by three different annotators. Annotators are required to include emotional descriptions in the motion annotation if they correspond to features in the book. If hands are predominant in the motion sequence, the body description will also emphasize their movements. For hand motions, three other annotators individually annotate finger-level actions for the left and right hand, focusing on the changing process of finger movements and gaining detailed records for prominent finger gestures. Fig. 2 offers a comprehensive exemplification of our dataset.
@@ 4 BOTH2Hands Algorithm
Based on our novel dataset, our objective is to generate hand motions that align with both textual prompts and body movements. To accomplish this, we propose a novel pipeline called BOTH2Hands, able to deal with any condition to generate lively two-hand motions. As shown in Fig. 3. Our framework consists of a two-stage mechanism: a pair of diffusion-based hand motion denoisers and a cross-attention structured transformer. In the first stage, we feed the body and text controls into two parallel diffusion models. In the second stage, we blend the hand motions generated by two-modality controls. In the case of body-conditioned hand diffusion, our goal is to generate hand motions that are coordinated with the body motions (Sec. 4.1). For text-conditioned hand diffusion, we first need to get local hands to make the denoise process more focused on gesture. However, these generated hands are not in the same coordinate system as the body-conditioned hands. To address this issue, we employ forward kinematics (FK), projecting the local hands into the body motion space, thereby avoiding gesture rotation errors while blending (Sec. 4.2). After that, we perform cross-attention motion blending between the text-conditioned hands and the body-conditioned hands (Sec. 4.3). This process ensures the generated hand motions effectively combine the dynamics of body motion with the explicit textual conditions.
@@ 4.1 Body-hand motion diffusion
Motion Diffusion Model. In our approach, we adopt the formulation suggested in the denoising diffusion probabilistic model (DDPM) [18], which enables us to effectively handle the hand synthesis task. The diffusion model processes the input data ($\textbf{x}_{0}$) by adding sampled Gaussian noise at first and then iteratively repeated for ${t}$ steps, resulting in the noised data at level ${t}$, denoted as $\textbf{x}_{t}$. This iterative process is commonly referred to as the forward process and can be represented as a Markov chain with $t$ steps, as shown in Eq. 1:
@@
where $\alpha_{t}=\prod_{i=1}^{t}{(1-\beta_{i})}$ and $\beta_{t}$ is a variance schedule parameters. The reverse diffusion process can be modeled as $p_{\theta}(x_{t-1}|x_{t},c_{0:N})$, the $\theta$ represents the learned parameters and $c_{0:N}$ represents a set of given conditions (00 indicates no condition). Notable, we can always train a diffusion denoiser with any condition, which can learn a Gaussian posterior distribution $q(x_{t-1}|x_{t},x_{0})$. The denoising sampling process can be formulated as:
@@
The term $\mu_{\theta}({x}_{t},t,c_{0:N})$ is the mean to learn, which can be influenced by the conditions $c_{0:N}$. As mentioned in previous work [10, 20], to be more exact, the updating rule of the mean is:
@@
Body to Hand Diffusion. As the same body motion may lead to different hand gestures, we need the diffusion probabilistic model to sample the most possible gestures instead of directly matching one gesture on the body. We utilize the widely adopted model SMPLH (SMPL+MANO) [31, 47] as our skeleton, with a total of $52$ joints, where the initial $22$ joints are body joints and the remaining $30$ joints are hand joints. We parameterize the representation of motions as positions and rotations 6D of joints. To fetch parameters, we use FK to calculate the absolute rotations and joint positions (we define absolute parameters as real positions and rotations of joints without the intervention of parent joints), denoted as global motions, where the former $22$ joints position and 6D corresponds to global body ($\textbf{C}_{B}\in\mathbb{R}^{T\times 22\times 9}$) and the latter $30$ joints corresponds to global hands ($\textbf{X}_{B}\in\mathbb{R}^{T\times 30\times 9}$). We do the forward process by adding noise to the hand motions step-by-step, fetching the sequences of hand motions $\textbf{X}^{t}_{1},\textbf{X}^{t}_{2},...,\textbf{X}^{t}_{T}$ at noise level $t$. Followed by the forward process, we conduct a reverse diffusion procedure on the transformer self-attention denoiser to estimate $\textbf{X}^{0}_{B}$. Then we adopt methods in EgoEgo [22] by directly concatenating on the noised hand $\textbf{X}^{t}_{B}$ and cleaned body $\textbf{C}^{0}_{B}$ as denoiser input, doing body-hand diffusion process, with the loss shown in Eq. 5:
@@ 4.2 Text-hand motion diffusion
For text-conditioned hand synthesis, we extract the local position from mocap data and turn them into rotation 6D. Then we discard body rotation, keeping hand rotation 6D as ground truth, the calculation process is defined below:
@@
where M is the full body motion aligned with text condition, $IK(\cdot)$ is inverse kinematics process and $Cat(\cdot,\cdot)$ is concatenate operation. For joint positions, we first use FK to calculate the $52$ absolute joint positions, then we can obtain hand joints in the origin of the coordinate system by subtracting the positions of their respective wrist for each hand joint:
@@
Hand Projection. Relative position (we define relative parameters as joint positions and rotations relative to parent joints) representation may result in motion drifting due to the need for integrating velocity to obtain absolute positions [57]. Nevertheless, relative rotation representations are advantageous for focusing on gestures and are easy to migrate [31]. Remember that we use motion representation, consisting of positions and rotations 6D. For body-conditioned hand synthesis, absolute positions and rotations of hand joints with body joints are directly applicable. However, predicting the absolute pose of hands without wrist positions is hard and meaningless for text-conditioned synthesis, so for the position, we prefer absolute representation in the origin of the coordinate system to avoid integration prediction and parent skeleton influence. To focus on the gesture itself, we preferred to use relative rotation. Then we got hand representations for text condition diffusion to train. Nonetheless, this representation cannot be used for hand blending directly, since the two conditioned hands are on different coordinate systems. In order to mitigate the influence of the spatial reference system, we projected $\textbf{x}_{T}$ in local space to $\textbf{X}_{T}$ in global space to eliminate their potential space error:
@@
Text to Hand diffusion. On the text-conditioned diffusion process, we follow the methods proposed in MDM [56] to add the text condition token c to embed noise $t$ step token, the denoising structure is similar to the body-conditioned motion denoiser, with a slightly different on input dimension, we only feed noised hand as input cause text-conditioned synthesis can not contain body motion. The loss type is reconstruction loss similar to Eq. 5 predicting $\textbf{x}^{0}_{T}$:

### 09 Apr 2024 v4

@@ Abstract
The recently emerging text-to-motion advances have inspired numerous attempts for convenient and interactive human motion generation. Yet, existing methods are largely limited to generating body motions only without considering the rich two-hand motions, let alone handling various conditions like body dynamics or texts. To break the data bottleneck, we propose BOTH57M, a novel multi-modal dataset for two-hand motion generation. Our dataset includes accurate motion tracking for the human body and hands and provides pair-wised finger-level hand annotations and body descriptions. We further provide a strong baseline method, BOTH2Hands, for the novel task: generating vivid two-hand motions from both implicit body dynamics and explicit text prompts. We first warm up two parallel body-to-hand and text-to-hand diffusion models and then utilize the cross-attention transformer for motion blending. Extensive experiments and cross-validations demonstrate the effectiveness of our approach and dataset for generating convincing two-hand motions from the hybrid body-and-textual conditions. Our dataset and code will be released to the community for future research, which can be found at github.
{strip}
@@ 1 Introduction
The recent text-to-motion advances [62, 29, 8, 74] mostly focus on generating body motions only. In contrast, the convenient generation of two-hand motions from text prompts has significantly fallen behind, mainly due to severe data scarcity. The wider adopted text-motion datasets [38, 14] embrace limited two-hand motions and corresponding textual annotations. Only recently, the concurrent work Motion-X [30] provides a large-scale dataset with expressive human motions and paired text prompts. Yet, it still lacks detailed annotations for the hand motions, making the fine-grained generation challenging, let alone enabling explicit finger-level controls. On the other hand, various methods [43, 51, 50] synthesize two-hand motions with body motion as extra conditions. Such a body-to-hand setting implicitly reasons the inherent correlations of human motions between the body and hands, and hence effectively handles specific scenarios like speeches or daily conversations [79, 71]. However, only body-level reasoning falls short of providing explicit and direct controls of hand motions, especially in a human-interpretable manner like text prompts.
@@
Table 1: Dataset comparisons. We conduct a comparison of datasets that encompass body and hand motions. Vocab. denotes the distinct vocabulary numbers used for annotation. Annotation refers to text annotations. Hand Joint Standard Deviation reflects the standard deviation of hand joint positions, indicating the diversity of hand motions. Detailed annotation refers to the number of text annotations for specific skeleton parts in each motion clip.
@@ 2 Related Works
Motion Generation. Currently, numerous works focus on motion generation under various conditions such as text and label [62, 74, 75, 47, 46, 14, 9, 1, 76, 44, 3], speech and music  [70, 34, 79, 69, 15, 78, 68] and objects [59, 4, 7, 16]. Other interesting works use brand-new algorithms [18] or focus on new scenes [31]. Among these, text-to-motion generation is a challenging task due to the difficulty of aligning natural language with time and space [45, 24]. MotionClip [61] aligns text with other modalities, enhancing model mapping text to motion. As diffusion model [56, 20, 57, 10] was introduced in various tasks [52, 55, 8, 52, 54], it also performs well in motion generation. For instance, Human Motion Diffusion Model [62] introduced text conditions and showed good results. Other works like T2M-GPT [74] applied the transformer architecture in this task and proved its effectiveness. MLD (motion-latent-diffusion) [6] has attempted to generate motions in the latent space. Some other works like InterGen [29] focus on human interaction scenes achieving good results. Full-body motion generation holds considerable significance in some specific domains like human object interaction (HOI) [59, 13, 64, 27, 60] and speech [70, 2, 35, 67], due to its extensive applicability.
@@
Hands are equally important as bodies in motion [42], presenting unique challenges due to their high density in small spatial occupancy. Previous hand generation works concentrate on physics-based issues [48, 32, 77] such as surface contact [72] and reconstructing hands [53, 28]. Other data-driven [22, 39, 41, 58] works often focus on specific scenarios [73], modeling the individual as a whole rather than considering parts separately [12, 65]. Some works like Body2Hands [43] take body as condition and achieve impressive results [51, 50, 71, 79]. Previous studies often focus on full-body or body motions only, rather than generating hands aligned with both body motions and text controls.
@@
Motion Dataset. These days various motion datasets have been presented. Action-labeled datasets like BABEL [49] offer verb-object phrases as conditions, which is unnatural for human communication. Datasets such as KIT [46] or HumanML3D [14] provide detailed natural annotations, while they ignore hands. Other datasets focus on hand scenarios like Hand-Object Manipulation [11], 3D Interacting Hand [26, 40]. Yet, such scenes mostly focus on hands, they hardly contain both hand and body data. Full-body datasets like GRAB [59] contain rich hand gestures but are narrowed down to HOI scenes. BEAT [33] uses speech text as conditions, lacking standard motion description. Currently, the largest full-motion dataset Motion-X [30] contains descriptions aligned with motions but lacks annotations focusing on hands, and the diversity of hand movement is less rich than their body motions. More data including rich daily hand gestures with detailed annotation is needed for body hand motion synthesis.
@@ 3 BOTH57M Dataset
Overview. We introduce the BOTH57M, a unique body-hand motion dataset comprising 1,384 motion clips and 57.4M frames, with 23,477 manually annotated motions and a rich vocabulary of 4,140 words. The dataset focuses on hands and body motion in daily various activities, referencing the book “Dictionary of Gestures ” [5] and supplementing with custom-designed movements. To the best of our knowledge, this is the only dataset that provides hybrid and detailed annotations of both body and hands at present, providing huge potential for future research. Tab. 1 shows a detailed comparison of various body hand datasets with ours. The rich vocabulary and hand diversity underscores our advantage in tackling the text/body-to-hand task.
Data Collecting. We utilize 32 RGB cameras to build a dense-view system for body-hand motion capturing. During data collection, participants are instructed to perform movements listed in the “Dictionary of Gestures” excluding unfriendly gestures. Subsequently, manual annotations are implemented. Three annotators are required to annotate full-body motions. For hand motions, three other annotators individually annotate finger-level actions for the left and right hand, focusing on the changing process of finger movements and gaining detailed records for prominent finger gestures. Fig. 2 offers a comprehensive exemplification of our dataset. For a comprehensive understanding of data collection and processing, as well as an in-depth explanation of our collected motions and text annotations, please refer to the supplementary material.
@@ 4 BOTH2Hands Algorithm
Based on our novel dataset, our objective is to generate hand motions that align with both textual prompts and body movements. To accomplish this, we propose a novel pipeline called BOTH2Hands to deal with rich conditions to generate lively two-hand motions, as shown in Fig. 3. Our framework consists of a two-stage mechanism: a pair of diffusion-based hand motion denoisers and a cross-attention structured transformer. In the first stage, we feed the body and text controls into two parallel diffusion models. In the second stage, we blend the hand motions generated by two-modality controls. Specifically, we follow EgoEgo [25] and adopt forward kinematics (FK) to get joint positions and rotations 6D in body motion space. For body-conditioned hand diffusion, our goal is to generate hand motions that are coordinated with the body motions (Sec. 4.1). As for text-conditioned hand diffusion, we first use inverse kinematics (IK) to get local hands to make the denoise process more focused on gesture. However, these generated hands are not in the same coordinate system as the body-conditioned hands. To address this issue, we then employ FK to project the local hands back into the body motion space, thereby eliminating gesture rotation errors while blending (Sec. 4.2). After that, we perform cross-attention motion blending between the text-conditioned hands and the body-conditioned hands (Sec. 4.3). This process ensures the generated hand motions effectively combine the dynamics of body motion with the explicit textual conditions.
@@ 4.1 Body-hand motion diffusion
Motion Diffusion Model. In our approach, we adopt the formulation suggested in the denoising diffusion probabilistic model (DDPM) [20], which effectively handles the hand synthesis task. The diffusion model processes the input data ($\textbf{x}_{0}$) for $t$ iterations and obtains the noised data at level $t$. In each iteration, sampled Gaussian noise is added to the data from the previous level. This iterative process is commonly referred to as the forward process and can be represented as a Markov chain with $t$ steps. The transition probability is shown in Eq. 1:
@@
where $\beta_{t}$ is a variance schedule parameter and $\alpha_{t}=\prod_{i=1}^{t}{(1-\beta_{i})}$. The reverse diffusion process can be modeled as $p_{\theta}(x_{t-1}|x_{t},c_{0:N})$, where $\theta$ represents the learned parameters and $c_{0:N}$ represents a set of given conditions (00 indicates no condition). Notable, we can always train a diffusion denoiser with any condition to learn a Gaussian posterior distribution $q(x_{t-1}|x_{t},x_{0})$. The denoising sampling process can be formulated as:
@@
The term $\mu_{\theta}({x}_{t},t,c_{0:N})$ is the mean to learn, which can be impacted by the conditions $c_{0:N}$. As mentioned in previous works [10, 23], to be more exact, the updating rule of the mean is:
@@
Body to Hand Diffusion. As the same body motion may lead to different hand gestures, we need the diffusion probabilistic model to sample the most possible gesture instead of directly matching one gesture on the body. We utilize the widely adopted model SMPLH (SMPL+MANO) [36, 53] as our skeleton, with a total of $52$ joints, where the initial $22$ joints are body joints and the remaining $30$ joints are hand joints. We parameterize the representation of motions as positions and rotations 6D of joint. To fetch parameters, we use FK to calculate the absolute rotations and joint positions (we define absolute parameters as real positions and rotations of joints without the intervention of parent joints), denoted as global motions, where the former $22$ joint positions and rotations 6D correspond to global body ($\textbf{C}_{B}\in\mathbb{R}^{T\times 22\times 9}$) and the latter $30$ joints correspond to global hands ($\textbf{X}_{B}\in\mathbb{R}^{T\times 30\times 9}$). We perform the forward process by adding noise to the hand motions step-by-step, fetching the sequences of hand motions $\textbf{X}^{t}_{1},\textbf{X}^{t}_{2},...,\textbf{X}^{t}_{T}$ at noise level $t$. Followed by the forward process, we conduct a reverse diffusion procedure on the transformer self-attention denoiser to estimate $\textbf{X}^{0}_{B}$. Then we adopt methods in [25] to directly concatenate the noised hand $\textbf{X}^{t}_{B}$ and cleaned body $\textbf{C}^{0}_{B}$ together as denoiser input during the body-hand diffusion process, with the loss shown in Eq. 5:
@@ 4.2 Text-hand motion diffusion
For text-conditioned hand synthesis, we use IK to extract the local positions and rotations 6D from FK motion results. Then we discard body rotations, keeping hand rotations 6D as ground truth, the calculation process is defined below:
@@
where M is the full body motion aligned with text condition, $IK(\cdot)$ is inverse kinematics process and $Cat(\cdot,\cdot)$ is concatenate operation. For joint positions, we first use FK to calculate the $52$ absolute joint positions. Then we can obtain hand joints in the origin of the coordinate system by subtracting the positions of their respective wrist for each hand joint:
@@
Hand Projection. Relative positions (we define relative parameters as joint positions and rotations relative to parent joints) representation may result in motion drifting due to the need for integrating velocity to obtain absolute positions [63]. Nevertheless, relative rotation representations are advantageous for focusing on gestures and are easy to migrate [36]. Remember that we use motion representation, consisting of positions and rotations 6D. For body-conditioned hand synthesis, absolute positions and rotations of hand joints with body joints are directly applicable. However, predicting the absolute pose of hands without wrist positions is hard and meaningless for text-conditioned synthesis, so for the positions, we prefer absolute representation in the origin of the coordinate system to avoid integration prediction and parent skeleton influence. To focus on the gesture itself, we prefer to use relative rotations. Nonetheless, absolute positions and relative rotations cannot be used for hand blending directly, since the two conditioned hands are on different coordinate systems. In order to mitigate the influence of the spatial reference system, we projected $\textbf{x}_{T}$ in local space to $\textbf{X}_{T}$ in global space to eliminate their potential space error:
@@
Text to Hand diffusion. On the text-conditioned diffusion process, we follow the methods proposed in MDM [62] to add the text condition token c to embed noise $t$ step token. The denoising structure is similar to the body-conditioned motion denoiser, with a slight difference in input dimension. We only feed the noised hands as input since text-conditioned synthesis can not contain body motion. We adopt the reconstruction loss similar to Eq. 5 to predict $\textbf{x}^{0}_{T}$: